# Unveiling Factual Recall Behaviors of Large Language Models through Knowledge Neurons

### v2

@@ 7.2 Results Analysis
The presence of knowledge conflict within the context consistently augments the faithfulness of LLMs in the corresponding fact. According to Figure 5 and Figure 6, the context of knowledge conflict results in the highest KN Scores of the corresponding hop fact, which indicates counterfactual context significantly improves the internal retrieval of that corresponding hop fact. It illustrates LLMs exhibit greater confidence in their encoded knowledge when confronted with knowledge conflict, a finding that aligns with the studies conducted by Zhou et al. (2023) and Li et al. (2023). When the knowledge presented in the context conflicts with the second-hop fact, it not only reinforces the retrieval of the second-hop fact but also enhances the recall of the first-hop fact. It is plausible that the introduction of the subject $o_{1}$ encourages LLMs to recall the precise triplet $(s,r_{1},o_{1})$. However, this effect does not extend to the first-hop fact. The occurrence of knowledge distraction appears not to cause much obstruction to the factual recall within LLMs. On the contrary, it may even stimulate LLMs to retrieve more facts sometimes, as evidenced by the high KN Scores for the first-hop fact of LLaMA2-7B when the knowledge distractor corresponding to the second-hop fact appears in Figure 6.
@@ A.1 Sampling two-hop factual triples
Our dataset is constructed based on Wikidata Vrandečić and Krötzsch (2014), a structurally optimized database covering nearly all domains. First we show manually selected relations that are used to construct two-hop relations:
@@ Appendix B Knowledge Neurons
Equation 10 is applied to the calculation of attribution scores for single-token target $o$. The method for computing attribution scores for multi-token target $o$ is described in Equation 11. Assuming the tokenized sequence of a relational-fact query and the corresponding ground truth respectively are $\left[q_{1},q_{2},\cdots,q_{n}\right]$ and $\left[gt_{1},gt_{2},\cdots,gt_{m}\right]$
@@
Appendix C Peformance of LLMs under CoT
C.1 CoT Results
Figure 9: The reasoning accuracy of LLaMA2-7B across different relations under three CoT settings
Figure 10: The reasoning accuracy of LLaMA3-8B across different relations under three CoT settings.
Figure 11: Distribution of intermediate layers in all FFNs for Mistral-7B.
Figure 12: Distribution of intermediate layers in all FFNs for LLaMA2-7B
Figure 13: Distribution of intermediate layers in all FFNs for LLaMA3-8B.
We present the reasoning accuracy of two additional models, LLaMA2-7B and LLaMA3-8B, on the filtered TFRKN dataset across various relations under three distinct conditions: no CoT, zero-shot CoT, and few-shot CoT. The comparative results are respectively illustrated in Figures 9 and 10. Figure 9 reveals that LLaMA2-7B faces significant challenges in addressing two-hop factual reasoning questions, particularly in comparison to LLaMA3-8B [^1] https://huggingface.co/meta-llama/Meta-Llama-3-8B-Instruct and Mistral-7B.
Further, Figures 11, 12, and 13 visually depict the activation patterns across the entire neuron spectrum of all intermediate layers within Feedforward Neural Networks (FFNs). Given the extensive neuron count, the heatmaps have been appropriately scaled for clarity. The vertical axis corresponds to the layer depth, while the horizontal axis indicates the neuron index within an intermediate layer. These visual representations effectively highlight the distribution of Knowledge Nodes (KNs), which are predominantly concentrated in the uppermost layers of LLMs.
Additionally, it is observable that the application of CoT activates a larger set of knowledge neurons during the factual reasoning process, thereby facilitating the LLMs’ ability to recall and extract internal knowledge. This observation reaffirms the pivotal role of CoT in the factual recall mechanism.
C.2 CoT prompting templates
In our experimental design, we have established two distinct prompting configurations for CoT: zero-shot CoT and few-shot CoT. In the case of zero-shot CoT, we simply precede the reasoning question with the directive "Let’s think step by step". Conversely, for few-shot CoT, we provide n-shot examples to guide the model’s reasoning process as in Figure 14.
Figure 14: The prompt of few-shot CoT.
Appendix D Experimental Details
@@
We present a comprehensive overview of our experimental setup. Our experiments are conducted using a refined subset of TFRKN dataset. To ensure that LLMs know each factual element required by the factual reasoning questions, we meticulously filtered out unqualified data points for each model. By taking the intersection of these filtered datasets, we culled a dataset comprising 1072 qualified data points. The process of identifying KNs for each fact triplet proves to be the most computationally intensive, with each model taking 96 GPU hours to find all KNs. In the context of the location experiment, we configured the integrated gradient steps to 20 and set the parameter of the shared percentage of coarse neurons to 0.2. The experiments were executed on a system equipped with NVIDIA A100 80GB GPUs, and further details of the software environment are available in our code repository. For knowledge conflict experiments, we construct a knowledge distraction sentence pool, randomly assigned to each reasoning question while knowledge conflict in cloze task is constructed by a set predefined templates of relations. All experimental results are the mean values of three repetitive experiments.

### v3

@@ 7.2 Results Analysis
The presence of knowledge conflict within the context consistently augments the faithfulness of LLMs in the corresponding fact. According to Figure 5 and Figure 6, the context of knowledge conflict results in the highest KN Scores of the corresponding hop fact  [^1] The results were obtained from a one-tailed paired sample t-test, conducted at a significance level of 0.05., which indicates counterfactual context significantly improves the internal retrieval of that corresponding hop fact. It illustrates LLMs exhibit greater confidence in their encoded knowledge when confronted with knowledge conflict, a finding that aligns with the studies conducted by Zhou et al. (2023) and Li et al. (2023). When the knowledge presented in the context conflicts with the second-hop fact, it not only reinforces the retrieval of the second-hop fact but also enhances the recall of the first-hop fact. It is plausible that the introduction of the subject $o_{1}$ encourages LLMs to recall the precise triplet $(s,r_{1},o_{1})$. However, this effect does not extend to the first-hop fact. The occurrence of knowledge distraction appears not to cause much obstruction to the factual recall within LLMs. On the contrary, it may even stimulate LLMs to retrieve more facts sometimes, as evidenced by the high KN Scores for the first-hop fact of LLaMA2-7B when the knowledge distractor corresponding to the second-hop fact appears in Figure 6.
@@ 9 Conclusions
Acknowledgements
This work was supported in part by the Strategic Priority Research Program of Chinese Academy of Sciences under Grant #XDA27030100 and the National Natural Science Foundation of China under Grants #72293573.
@@ A.1 Sampling two-hop factual triples
Our dataset is constructed based on Wikidata Vrandečić and Krötzsch (2014), a structurally optimized database covering nearly all domains. The dataset is available at https://github.com/wangyifei0047/TFRKN.
First we show manually selected relations that are used to construct two-hop relations:
@@ Appendix B Knowledge Neurons
Equation 10 is applied to the calculation of attribution scores for single-token target $o$. The method for computing attribution scores for multi-token target $o$ is described in Equation 11. Assuming the tokenized sequence of a relational-fact query and the corresponding ground truth respectively are $\left[q_{1},q_{2},\cdots,q_{n}\right]$ and $\left[gt_{1},gt_{2},\cdots,gt_{m}\right]$.
@@
Appendix C Experimental Details
We present a comprehensive overview of our experimental setup.
Intersection of LLMs
Experiments are conducted using a refined subset of TFRKN dataset. To ensure that LLMs know each factual element required by the factual reasoning questions, we meticulously filtered out unqualified data points for each model. By taking the intersection of these filtered datasets, we culled a dataset comprising 1072 qualified data points.
Indentification of KNs
@@ Appendix C Experimental Details
The process of identifying KNs for each fact triplet proves to be the most computationally intensive, with each model taking 96 GPU hours to find all KNs. In the context of the location experiment, we configured the integrated gradient steps to 20 and set the parameter of the shared percentage of coarse neurons to 0.2. The experiments were executed on a system equipped with NVIDIA A100 80GB GPUs, and further details of the software environment are available in our code repository. All experimental results are the mean values of three repetitive experiments.
Appendix D Construction of Contextual Conflict
Knowledge Distraction
We manually constructed a set of irrelevant fact statements $S$. $S$ does not involve any entities or relations in TFRKN to ensure "unrelated" property. Each two-hop question randomly selects a knowledge distraction from this set.
Knowledge Conflict
We constructed contexts that conflict with the first-hop fact and that conflict with the second-hop fact for each two-hop question respectively. The method is as follows: we manually designed templates $T$ for all relations involved in the TFRKN dataset. Assuming there is a fact $(s,r,o)$, we collect the set of candidate objects related to $r$ in the dataset, select an $o^{*}$ that is not equal to $o$ as the new fabricated fact $(s,r,o^{*})$, and apply the template of the relation in $T$ to obtain the knowledge conflict context corresponding to $(s,r,o)$.
Appendix E Additional Experimental Results
LLaMA2 LLaMA3 Mistral
Avg. Med. Max. Avg. Med. Max. Avg. Med. Max.
Number of KNs per fact 26.4 22.0 53.0 28.3 25.0 52.0 22.3 26.0 59.0
Number of pairwise intersections 2.7 2.0 6.0 3.7 3.0 4.0 1.0 0.0 4.0
Table 7: The KNs for different facts may vary significantly (Avg.: average, Med.: median, Max.: maximum).
Sentences ending with Paris LLaMA2-7B LLaMA3-8B Mistral-7B
The capital of France is 1.98 2.25 2.04
The capital city of France is 2.01 2.27 2.06
The Louvre Museum is situated in the city of 1.10 1.46 1.55
The Seine River flows gracefully through the heart of 1.27 1.47 1.70
The City of Love refers to the city of 1.24 1.70 1.41
The City of Light refers to the city of 1.25 1.65 1.53
The capital The football club Paris Saint-Germain is based in the city of France is 1.24 1.40 1.66
The Eiffel Tower is one of the most iconic landmarks in the city of 1.37 1.80 1.64
Table 8: KN Scores corresponding to $(France,capital,Paris)$ for different sentences which end with Paris.
Non-overlap of Knowledge Neurons
Based on the KNs identified in our experiments, we conducted a verification of non-overlap. To achieve this, we randomly sampled 6,000 pairs of distinct relational facts and calculated the number of intersecting KNs between each pair. The statistics of overlapping KNs are shown in Table 7.
Verification of Basic Assumptions
We present compelling results from small-scale case studies, which prove that when LLMs predict the same word, the metric of KN Scores would be high only when the process involves fact retrieval. For illustration, we use $(France,capital,Paris)$ as an example, whose KNs cover 26 neurons. KN Scores are computed across these 26 neurons as the metric. Then we construct sentences that end with "Paris" and then replace "Paris" with a blank, prompting LLMs to predict the missing word. To ensure that the LLMs predict "Paris" as the final token, we design straightforward and commonsense sentences and verify that the LLMs would indeed predict "Paris" and then assess the knowledge-expressing prompts and compare them with non-knowledge-expressing prompts by analyzing their KN Scores.
"The capital of France is" and "The capital city of France is" are knowledge-expressing prompts, which consistently exhibit higher KN Scores compared to other examples, even though LLMs predict "Paris" for all these sentences. This experiment illustrates that KNs for $(France,capital,Paris)$ are activated mostly when LLMs recall $(France,capital,Paris)$, not when they make a specific predictive word "Paris".